OpenStudy (anonymous):

help please: Multiply and simplify if possible 4^sqrt11 * 4^sqrt10 a. 11^4sqrt10 b. 11 c. 10 d. 4^sqrt110

OpenStudy (jhannybean):

Remember: \(\sqrt{m} \cdot \sqrt{n} = \sqrt{ m\cdot n}\)

OpenStudy (jhannybean):

And I'm not quiet understanding what your notation means. Would you kindly draw your problem using the drawing tool,or post a printscreen of your problem? :)

OpenStudy (jhannybean):

I'll take a guess:\[\large \sqrt[4]{11} \cdot \sqrt[4]{10}\] Would that be your problem?

OpenStudy (anonymous):



OpenStudy (anonymous):

exACTLY.

OpenStudy (jhannybean):

Oh ok, got it :)

OpenStudy (jhannybean):

So from my first post, when you multiply 2 numbers under a square root, you can easily combine them into 1 square root by multiplying the numbers together.. What is \(11 \cdot 10 =~?\)

OpenStudy (anonymous):

110